# Thorough Analysis of $\pi^{-}p\rightarrow K^{*}\Sigma$ Reaction: Probing Nucleon Excited State and Constituent Scaling Laws

### v1

@@ III Numerical results
III.1 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$
Using the MINUIT code from CERNLIB, we will fit the experimental data CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Both total and differential cross section datas are used in a $\chi^{2}$ fitting algorithm to determine the values of the free parameters. The fitting parameters are listed in Table 1. The fit excluding the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ yields $\chi^{2}/d.o.f.$ = 6.19, while including the $N^{*}(2080)(\frac{3}{2}^{-})$ improves the fit to $\chi^{2}/d.o.f.$ = 1.52. This indicates a significant contribution from the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$.
Table 1: Fitted values of the free parameters with all experimental data in Ref. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$.
with $N^{*}(2080)(\frac{3}{2}^{-})$ without $N^{*}(2080)(\frac{3}{2}^{-})$
@@ III.1 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$
The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ having $J^{P}$ = $\frac{3}{2}^{-}$ is illustrated in Fig. 2. The results, which include the contributions from the $t$-channel, $u$-channel, and $s$-channel, show good agreement with the experimental data. Additionally, Fig. 2 demonstrates that while the contributions from the $t$- and $u$-channels are relatively small, the $s$-channel plays a crucial role in the reaction.
Figure 2: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 1. The solid (red), dashed-dotted (blue), dashed (purple), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{3}{2}^{-}$.
@@
In Fig. 3, we present the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ as a function of $\cos\theta$, specifically for the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$ at a center-of-mass energy $W=3.056$ GeV. The analysis reveals that the $t$-channel contribution dominates at forward angles, while the $u$-channel contribution is small and negligible. In Fig. 4, we show the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of center-of-mass energy, with $\cos\theta=1$, for the same $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$. Since these measurements are taken at forward angles, the distinct shapes of the curves resulting from the contributions of the three channels allow us to effectively isolate the contributions from the $t$-channel and $u$-channel, while providing a clear indication of the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$. This approach is ideal for studying the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ through differential cross sections at forward angles.
In Fig. 5, we calculate the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. The results indicate that as the energy approaches the threshold of $K^{*}\Sigma$, the contribution of the $s$-channel becomes more prominent. Conversely, as energy increases, the contribution of the $t$-channel at forward angles becomes more significant, while the $u$-channel contribution remains consistently small. These findings suggest that measuring the differential cross section of the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ near the $K^{*}\Sigma$ threshold will provide valuable information about the $N^{*}(2080)(\frac{3}{2}^{-})$ resonance.
@@
The $t$-distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ and $J^{P}=\frac{3}{2}^{-}$ at $W=2.88$ GeV is shown in Fig. 6. The results, incorporating full contributions from the $s$-channel, $t$-channel, and $u$-channel, are in good agreement with the experimental data. Fig. 7 presents the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. These results provide insight into the production mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, aiding in the differentiation of its underlying processes.
@@
III.2 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$
Similar to the analysis of $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, we also calculate the cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$ as detailed in Figs. 8, 9 and 10. We employ the same fitting scheme as in Table 1, treating the cutoff value as a free parameter. The fitting parameters are shown in Table 2. The fit excluding the $N^{*}(2080)(\frac{1}{2}^{-})$ contribution yields a $\chi^{2}/d.o.f.$ = 6.19, whereas including this contribution results in $\chi^{2}/d.o.f.$ = 2.20. For the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, the $\chi^{2}/d.o.f.$ = 1.52, indicating a better fit to the experimental data for the $N^{*}(2080)(\frac{3}{2}^{-})$.
Fig. 9 and Fig. 10 present the differential cross section $d\sigma/d\cos\theta$ and the $t$-distribution for $\pi^{-}p\rightarrow K^{*}\Sigma$, respectively. These results closely match those in Fig. 3 and Fig. 6. Fig. 8 shows the total cross section for $\pi^{-}p\rightarrow K^{*}\Sigma$, with a peak value approximately 500 $\mu b$ higher than in Fig. 2. In Fig. 11, the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of c.m. energies at $\cos\theta$ = 1 is presented. This provides an ideal means to assess the contribution of the $N^{*}(2080)(\frac{1}{2}^{-})$ through forward-angle differential cross sections. Fig. 12 shows the differential cross section at various c.m. energies. As the energy approaches the $K^{*}\Sigma$ threshold, the $s$-channel contribution becomes more pronounced. With increasing energy, the $t$-channel contribution at forward angles grows more significant, while the $u$-channel contribution remains small. This behavior is similar to that observed for the $N^{*}(2080)(\frac{3}{2}^{-})$. Fig. 13 displays the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies. These results may aid us in distinguishing the production mechanisms for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction with the $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$.
Table 2: Fitted values of the free parameters with all experimental data in Refs. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{1}{2}^{-}$.
with $N^{*}(2080)(\frac{1}{2}^{-})$ without $N^{*}(2080)(\frac{1}{2}^{-})$
@@ III.2 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$
Figure 8: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 2. The solid (purple), dashed-dotted (blue), dashed (red), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{1}{2}^{-}$.
@@
IV constituent counting rule in hard exclusive reactions
We introduce the constituent counting rule for hard exclusive reactions by presenting the expressions for the cross section and matrix element, and by explaining a typical hard gluon exchange process that contributes to such reactions. For a large-angle exclusive scattering process $a+b\rightarrow c+d$, the reaction cross section is given by:
$\frac{d\sigma_{ab\to cd}}{dt}\simeq\frac{1}{16\pi s^{2}}\overline{\sum_{\mathrm{pol}}}|M_{ab\to cd}|^{2},$ (30)
In this context, the approximation implies that particle masses are neglected. The Mandelstam variables $s$ and $t$ are defined as follows:
[…]
where $p_{i}$ ($i=a,b,c,d$) represents the momentum of hadron $i$, while $p_{\mathrm{c.m.}}$ and $\theta_{\mathrm{c.m.}}$ denote the momentum and scattering angle in the center of mass (c.m.) frame, respectively. Eq. (30) includes an average over the initial spins and a sum over the final spins.
The matrix element $M_{ab\rightarrow cd}$ can be expressed in a factorized form when there is a large momentum transfer Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Efremov:1979qk ; Guo:2016fqg :
[…]
Here, $[x]$ represents the set of light-cone momentum fractions of partons within a hadron, where $x_{i}=p^{+}_{i}/p^{+}$, with $p_{i}$ and $p$ denoting the momenta of the $i$th parton and the hadron, respectively. The light-cone component is defined as $p^{+}=(p^{0}+p^{3})/\sqrt{2}$, considering the third axis as the longitudinal direction.
For a nucleon, two independent variables are required to describe the distribution amplitude, while accounting for the constraint of momentum conservation, $x_{1}+x_{2}+x_{3}=1$. Thus, we have $[x]=x_{1},x_{2}$ and $[dx]=dx_{1}dx_{2}$ in Eq. (32). Conversely, only one variable, $x$, is needed for mesons such as pions and kaons. The variable $Q^{2}$ represents the hard scale of the reaction, approximately given by $Q^{2}\simeq s$ for large-angle elastic exclusive scattering. In Eq. (32), the dependencies on the renormalization and factorization scales are suppressed: the former is governed by the renormalization group equation for the coupling constant, and the latter by the Efremov-Radyushkin–Brodsky–Lepage (ERBL) evolution equations for the distribution amplitudes Lepage:1979zb ; Efremov:1979qk . These scales are typically taken to be on the order of $Q^{2}$ to minimize large radiative corrections. Consequently, the $Q^{2}$ dependencies are logarithmic and do not significantly affect the scaling behavior of the matrix element.
Originally, the constituent counting rule was proposed based on dimensional counting Brodsky:1973kr ; Brodsky:1974vy ; Matveev:1973ra . It was subsequently explored within the framework of perturbative QCD, incorporating hard scattering processes Farrar:1979aw ; Efremov:1978rn ; Duncan:1979hi ; Duncan:1980qd ; Lepage:1980fj . In this section, we derive the scaling rule by counting mass dimensions.
Given that the state vector of a hadron is normalized as $\langle h(p^{\prime})|h(p)\rangle=2p^{0}(2\pi)^{3}\delta^{(3)}(\vec{p}-\vec{p}^{\prime})$, we find that its mass dimension is $[1/M]$. If a hadron is composed of $n_{l}$ elementary constituents, its state vector can be expressed as
$|l\rangle=\sqrt{N_{l}}|n_{l}\rangle,\quad\left[\sqrt{N_{l}}\right]=\left[M^{n_{l}-1}\right].$ (33)
The second equation indicates that the normalization factor $\sqrt{N_{l}}$ carries the mass dimension $\left[M^{n_{l}-1}\right]$ when the state vector of each constituent has the mass dimension $[1/M]$. In practice, this normalization factor corresponds to the “decay constant,” which is independent of the hard momentum scale Kawamura:2013iia .
We now consider the mass dimensions of the matrix element in Eq. (30) to derive the counting rule for exclusive cross sections. The scattering matrix $S$ is connected to the transition matrix $T$ by the relation $S=\bm{1}+i(2\pi)^{4}\delta^{(4)}(p_{f}-p_{i})T$, indicating that the mass dimension of $T$ is $[T]=[M^{4}]$. The matrix element $M_{ab\rightarrow cd}$ is then expressed in terms of $T$ as follows:
$M_{ab\to cd}=\langle cd|T|ab\rangle=\sqrt{N_{a}N_{b}N_{c}N_{d}}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle.$ (34)
Since the normalization factors $N_{i}(i=a,b,c,d)$ are represented by soft constants, we analyze the matrix element $\hat{M}_{ab\rightarrow cd}$ by omitting these factors. Consequently, the remaining hard component should be expressed in terms of the variables $s$ and $t$:
[…]
Given the dimensions $[T]=[M^{4}]$ and $[|n_{i}\rangle]=[1/M^{n_{i}}]$, the dimension of the matrix element is given by
$\begin{bmatrix}\hat{M}_{ab\to cd}\end{bmatrix}=\begin{bmatrix}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle\end{bmatrix}=\begin{bmatrix}M^{4-n}\end{bmatrix},$ (36)
where $n\equiv n_{a}+n_{b}+n_{c}+n_{d}$. In a large-angle exclusive reaction, the variable $s$ can be chosen as the sole hard scale. Thus, considering the mass dimension, the matrix element is expressed as:
$\hat{M}_{ab\to cd}=\hat{F}_{ab\to cd}(s,t)=s^{(4-n)/2}F_{ab\to cd}(t/s),$ (37)
where $F_{ab\to cd}(t/s)$ is a dimensionless function of scattering angle, with $-2t/s=1-\cos\theta_{c.m.}$ from Eq. (31). Using Eqs. (30), (35), and (37), we derive the constituent-counting expression for the cross section:
@@
where $f(t/s)$ accounts for the scattering angle and is multiplied by the normalization factors. The mass dimensions of $f(t/s)$ are given by $[f(t/s)]=[N_{a}N_{b}N_{c}N_{d}|F_{ab\to cd}|^{2}]=[M^{2n-8}]$, so the overall mass dimension of Eq. (38) is $[1/M^{4}]$. This derivation illustrates the counting rule based on mass dimensions, referred to as the “constituent counting rule” since it quantifies the number of constituents involved in the reaction.
Experimentally, the constituent counting rule was first applied in Ref. Anderson:1976ph to reactions such as $\gamma p\rightarrow\pi^{+}n$, $\pi^{0}p$, $\pi^{-}\Delta^{++}$, $\rho^{0}p$, $K^{+}\Lambda$, and $K^{+}\Sigma^{0}$, suggesting $n=3$ for baryons and $n=2$ for mesons. The JLab (Thomas Jefferson National Accelerator Facility) Hall-A and E94-104 collaborations subsequently verified that experimental data are consistent with $n_{\gamma}=1$, $n_{p}=n_{n}=3$, and $n_{\pi^{+}}=2$ by analyzing the $\gamma p\rightarrow\pi^{+}n$ reaction JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp . Furthermore, several experiments have confirmed the scaling behavior of the exclusive cross section predicted by the constituent counting rule Baller:1988tj ; White:1994tj .
Table 3: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$.
@@
Table 4: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$.
@@
In this study, we test the constituent counting rule using the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. According to the rule, the total number of constituents is $n=2+3+2+3=10$. We calculate the cross sections for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process at $\theta_{\mathrm{c.m.}}=90^{\circ}$, considering the cases where the $N^{*}(2080)$ has $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$. The results are presented in Table 3 and Table 4, respectively.
By fitting the experimental data at $\sqrt{s}=2.88$ GeV and the numerical results shown in Table 3 and Table 4 to the expression $d\sigma/dt=(\text{constant})\times s^{2-n}$, the scaling factors obtained are presented in Table 5. We find that for both $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, the fitted values of $n$ are very close to 10, confirming the accuracy of the constituent counting rule. Fig. 14 and Fig. 15 depict the differential cross section $d\sigma/dt$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at a large meson production angle ($\theta=90^{\circ}$ in the center of mass frame) as a function of $\sqrt{s}$ for the $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, respectively.
@@
In this work, the constituent counting rule of perturbative QCD could be utilized to determine the internal quark configuration of exotic-hadron candidates in exclusive production processesKawamura:2013iia . In the exclusive process $a+b\rightarrow c+d$ with large-momentum transfer, quarks should share large momenta so that they can stick together to form a hadron by exchanging hard gluons. Since the cross section of the $a+b\rightarrow c+d$ exclusive reaction should scale like $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$ with $n=n_{a}+n_{b}+n_{c}+n_{d}$, where $n_{i}(i=a,b,c,d)$ is the number of constituents in the particle $i$. Since the factor $n_{i}$ clearly indicates the internal configuration in the hadron, this scaling relation can be employed to find the internal configurations of exotic hadrons. The constituent counting rule provides useful guidance for us to further explore the properties of exotic hadrons.
V Dalitz process and experimental feasibility

### v2

@@ III Numerical results
III.1 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$
Using the MINUIT code from CERNLIB, we will fit the experimental data CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Both total and differential cross section datas are used in a $\chi^{2}$ fitting algorithm to determine the values of the free parameters. The fitting parameters are listed in Table 1. The fit excluding the contribution of the $P_{s\bar{s}}[3/2^{-}]$ yields $\chi^{2}/d.o.f.$ = 6.19, while including the $P_{s\bar{s}}[3/2^{-}]$ improves the fit to $\chi^{2}/d.o.f.$ = 1.52. This indicates a significant contribution from the $s$-channel $P_{s\bar{s}}[3/2^{-}]$.
Table 1: Fitted values of the free parameters with all experimental data in Ref. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$.
with $P_{s\bar{s}}[3/2^{-}]$ without $P_{s\bar{s}}[3/2^{-}]$
@@ III.1 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$
The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $P_{s\bar{s}}$ having $J^{P}$ = $\frac{3}{2}^{-}$ is illustrated in Fig. 2. The results, which include the contributions from the $t$-channel, $u$-channel, and $s$-channel, show good agreement with the experimental data. Additionally, Fig. 2 demonstrates that while the contributions from the $t$- and $u$-channels are relatively small, the $s$-channel plays a crucial role in the reaction.
Figure 2: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 1. The solid (red), dashed-dotted (blue), dashed (purple), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $P_{s\bar{s}}$ is $\frac{3}{2}^{-}$.
@@
In Fig. 3, we present the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ as a function of $\cos\theta$, specifically for the $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$ at a center-of-mass energy $W=3.056$ GeV. The analysis reveals that the $t$-channel contribution dominates at forward angles, while the $u$-channel contribution is small and negligible. In Fig. 4, we show the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of center-of-mass energy, with $\cos\theta=1$, for the same $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$. Since these measurements are taken at forward angles, the distinct shapes of the curves resulting from the contributions of the three channels allow us to effectively isolate the contributions from the $t$-channel and $u$-channel, while providing a clear indication of the $s$-channel $P_{s\bar{s}}[3/2^{-}]$. This approach is ideal for studying the contribution of the $P_{s\bar{s}}[3/2^{-}]$ through differential cross sections at forward angles.
In Fig. 5, we calculate the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. The results indicate that as the energy approaches the threshold of $K^{*}\Sigma$, the contribution of the $s$-channel becomes more prominent. Conversely, as energy increases, the contribution of the $t$-channel at forward angles becomes more significant, while the $u$-channel contribution remains consistently small. These findings suggest that measuring the differential cross section of the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ near the $K^{*}\Sigma$ threshold will provide valuable information about the $P_{s\bar{s}}[3/2^{-}]$ resonance.
@@
The $t$-distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $P_{s\bar{s}}$ and $J^{P}=\frac{3}{2}^{-}$ at $W=2.88$ GeV is shown in Fig. 6. The results, incorporating full contributions from the $s$-channel, $t$-channel, and $u$-channel, are in good agreement with the experimental data. Fig. 7 presents the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. These results provide insight into the production mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, aiding in the differentiation of its underlying processes.
@@
III.2 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{1}{2}^{-}$
Similar to the analysis of $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$, we also calculate the cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with $P_{s\bar{s}}$ having $J^{P}$ = $\frac{1}{2}^{-}$ as detailed in Figs. 8, 9 and 10. We employ the same fitting scheme as in Table 1, treating the cutoff value as a free parameter. The fitting parameters are shown in Table 2. The fit excluding the $P_{s\bar{s}}[1/2^{-}]$ contribution yields a $\chi^{2}/d.o.f.$ = 6.19, whereas including this contribution results in $\chi^{2}/d.o.f.$ = 2.20. For the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$, the $\chi^{2}/d.o.f.$ = 1.52, indicating a better fit to the experimental data for the $P_{s\bar{s}}[3/2^{-}]$.
Fig. 9 and Fig. 10 present the differential cross section $d\sigma/d\cos\theta$ and the $t$-distribution for $\pi^{-}p\rightarrow K^{*}\Sigma$, respectively. These results closely match those in Fig. 3 and Fig. 6. Fig. 8 shows the total cross section for $\pi^{-}p\rightarrow K^{*}\Sigma$, with a peak value approximately 500 $\mu b$ higher than in Fig. 2. In Fig. 11, the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of c.m. energies at $\cos\theta$ = 1 is presented. This provides an ideal means to assess the contribution of the $P_{s\bar{s}}[1/2^{-}]$ through forward-angle differential cross sections. Fig. 12 shows the differential cross section at various c.m. energies. As the energy approaches the $K^{*}\Sigma$ threshold, the $s$-channel contribution becomes more pronounced. With increasing energy, the $t$-channel contribution at forward angles grows more significant, while the $u$-channel contribution remains small. This behavior is similar to that observed for the $P_{s\bar{s}}[3/2^{-}]$. Fig. 13 displays the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies. These results may aid us in distinguishing the production mechanisms for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction with the $P_{s\bar{s}}$ having $J^{P}$ = $\frac{1}{2}^{-}$.
Table 2: Fitted values of the free parameters with all experimental data in Refs. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $P_{s\bar{s}}$ with $J^{P}=\frac{1}{2}^{-}$.
with $P_{s\bar{s}}[1/2^{-}]$ without $P_{s\bar{s}}[1/2^{-}]$
@@ III.2 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{1}{2}^{-}$
Figure 8: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 2. The solid (purple), dashed-dotted (blue), dashed (red), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $P_{s\bar{s}}$ is $\frac{1}{2}^{-}$.
@@
III.3 A further test by the constituent counting rule
In the following, we adopt the approach of the constituent counting rule to further test our theoretical calculation around the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Generally, for a large-angle exclusive scattering process $a+b\rightarrow c+d$, the reaction cross section can be expressed with the constituent counting rule (see Refs. Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Efremov:1979qk ; Brodsky:1973kr ; Brodsky:1974vy ; Matveev:1973ra ; Efremov:1978rn ; Duncan:1979hi ; Duncan:1980qd ; Lepage:1980fj ; Farrar:1979aw ; Kawamura:2013iia ; Anderson:1976ph ; JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp ; Baller:1988tj ; White:1994tj for more details):
@@ III.3 A further test by the constituent counting rule
where $n\equiv n_{a}+n_{b}+n_{c}+n_{d}$, while $f(t/s)$ accounts for the scattering angle and is multiplied by the normalization factors.
Table 3: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$.
@@
Table 4: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{1}{2}^{-}$.
@@
According to the rule, in the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, the total number of constituents is $n=2+3+2+3=10$. We calculate the cross sections for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process at $\theta_{\mathrm{c.m.}}=90^{\circ}$, considering the cases where the $P_{s\bar{s}}$ has $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$. The results are presented in Table 3 and Table 4, respectively. By fitting the experimental data at $\sqrt{s}=2.88$ GeV and the numerical results shown in Table 3 and Table 4 to the expression $d\sigma/dt=(\text{constant})\times s^{2-n}$, the scaling factors obtained are presented in Table 5. We find that the fitted values of $n$ are very close to 10, confirming the accuracy of the theoretical prediction for the cross section,. Fig. 14 and Fig. 15 depict the differential cross section $d\sigma/dt$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at a large meson production angle ($\theta=90^{\circ}$ in the center of mass frame) as a function of $\sqrt{s}$ for the $P_{s\bar{s}}[3/2^{-}]$ and $P_{s\bar{s}}[1/2^{-}]$, respectively. Both figures clearly demonstrate that as energy increases, the theoretically predicted differential cross section aligns more closely with the constituent counting rules. This occurs because these rules are generally more applicable in high-energy, large-momentum transfer regions. In this high-energy regime, which is distant from the mass threshold of $P_{s\bar{s}}$, the contribution of $P_{s\bar{s}}$ production to the differential cross section becomes minimal. Consequently, this region does not allow the constituent counting rules to reflect the properties of the intermediate exchange particle $P_{s\bar{s}}$. Introducing the constituent counting rules serves to verify the theoretical cross-section calculations presented in this work, further supporting the reliability of our theoretical model.
@@
IV Dalitz process and experimental feasibility